Ali deposited $1,200 at the end of every month into an RRSP for 6 years. The interest rate earned was 5.50% compounded semi-annually for the first 3 years and changed to 5.75% compounded monthly for the next 3 years. What was the accumulated value of the RRSP at the end of 6 years?
Answer #1

Effective rate = (1 + r / n)^n - 1
= (1 + 5.50% / 2)^2 - 1
= 5.5756%

Monthly rate = (1 + 5.5756%)^(1/12) - 1 = 0.4532%

Future value at the end of the 3 years can be calculated by using the following excel formula:
=FV(rate,nper,pmt,pv)
=FV(0.4532%,3*12,-1200,0)
= $46,808.68


Accumulated value at the end of 6 years can be calculated by using the following excel formula:
=FV(rate,nper,pmt,pv)
=FV(5.75%/12,36,-1200,-46808.68)
= $102,625.77

Accumulated value at the end of 6 years = $102,625.77